OpenStudy (anonymous):

assuming 100% dissociation, calculate the freezing point and boiling point of a 2.5 m SnCl4 (aq)

OpenStudy (anonymous):

100% dissociation means you have Sn^4+ + 4Cl^-, giving you an i (or van't Hoff factor) of 5. You're in an aqueous solution, so we know that water freezes at 0C, and boils at 100C. The equation you're looking for is: \[\Delta T = iK _{f}m\] Where i = van't hoff factor, Kf = freezing point depression constant, and m = molality. Kf is a constant, 1.86 °C kg/mol i is 5 as stated above molality.... 2.5 moles SnCL4 per 1 L water, want molality (mol solute/kg solvent) 1L water = 1000g at standard temp/pressure. 1.000kg 2.5 mol SnCl4/1.000kg H2O = 2.5m Temp change = 5 (unitless) * 25 mol/kg * 1.86 °C kg/mol Temp change = 23°C So the freezing point is now -23°C because adding solutes lowers the freezing point. Boiling point, same business except the constant changes, as does whether you add/substract your change in temperature. \[\Delta T = iK _{b}m\] Kb = 0.512 °C kg/mol Temp change = 5 * 2.5 mol/kg * 0.512 °C kg/mol = 6.4 °C Your new boiling point is 106°C. Viola! Hope this helped.